# Bayesian Invariance Modeling of Multi-Environment Data

### v1

@@ 1 Introduction
Originally, Peters et al. (2016) motivated these assumptions through a causal problem. Each environment arises by an unknown intervention on the features $x$; the outcome $y$ is directly caused by a subset of those features; and the goal is to analyze the multi-environment data to determine the direct causes of $y$. Here we solve the same problem but state the invariance assumption directly without appealing to causality. While these assumptions often arise from a causal problem, invariance has wider applications including predictive robustness (Bühlmann, 2020), domain adaptation (Rojas-Carulla et al., 2018), and out-of-distribution generalization (Arjovsky et al., 2019).
@@
The paper is organized as follows. In § 2 we state the main invariance assumption and introduce BIP, including the model structure, exact inference procedure, and the scalable variational approximation. In § 3, we present theoretical results on posterior consistency and contraction rates. In § 4, we validate BIP on simulations. In § 5, we use BIP to study real data. In § 6, we summarize BIP and discuss limitations and future work.
@@ 1.1 Related works
Since Peters et al. (2016), invariant prediction has become an active area of research. Heinze-Deml et al. (2018) extends this framework to non-linear settings. Rothenhäusler et al. (2019) proposes an invariance notion based on feature-residual inner products in linear models, inking it to causal inference under additive interventions. Fan et al. (2023) incorporates residual-based invariance as a regularization term in linear regression. Wang et al. (2024) establishes a theoretical connection between invariant prediction models and causal outcome models, leading to a computationally efficient algorithm for causal discovery. Gu et al. (2025) shows that solving for exact invariance is NP-hard and introduces an objective that interpolates between exact invariance and predictive performance.
This paper offers a new perspective on invariant prediction, by introducing a Bayesian approach. In this sense, our goals are most close aligned to Peters et al. (2016) and Fan et al. (2023). As we will see below, our probabilistic perspective on this problem opens the door to both new algorithms and new theoretical understanding.
@@ 2.1 Multi-environment data and the invariance assumption
We call $z^{*}$ the (true) invariant feature selector and $x^{z^{*}}$ the (true) invariant features. As a consequence of Assumption 1 and the chain rule, each environment’s joint distribution factorizes,
@@ 1 Initialize variational parameter $\phi_{0}$
In detail, we first estimate $\hat{g}(y\mid x^{z})$ and $\hat{p}_{e}(y\mid x^{z})$ from data, and then optimize Eq. 7 with stochastic optimization. Specifically, we use Monte Carlo approximations of the reparamaterization gradient, a form of black box variational inference (Kingma and Welling, 2014; Ranganath et al., 2014). Algorithm 2 summarizes this procedure. The complexity of VI-BIP is $O(T\cdot M\cdot c(\mathcal{D},p))$ for $T$ iterations and $M$ gradient samples. When the cost of exact BIP inference is prohibitive, VI-BIP provides a good alternative.
A technical challenge in VI-BIP is computing gradients of the ELBO, which involve expectations with respect to the discrete variational distribution $q_{\phi}(z)$. Standard reparameterization gradients do not directly apply, as they require continuous latent variables. Here we use the unbiased U2G gradient estimator (Yin et al., 2020); see Algorithm 3 in the supplementary materials. For a comprehensive discussion of gradient estimation methods for discrete latent variables, also see Mohamed et al. (2020). Additional implementation and optimization details for VI-BIP are provided in § B.2.
@@ 3 Theory
We use $\stackrel{{\scriptstyle P}}{{\to}}$ to denote convergence in probability, and $\forall_{p}z$ to mean for every $z$ with positive prior value. Additional results and proofs are included in Supplementary Section A.
@@ 3.1 Main results
The first is the prior, which influences the term $R$ in Eq. 8. An informative prior that assigns lower probability to non-invariant $z$s will lead to smaller value of $\max_{z\neq z^{*}}\frac{p(z)}{p(z^{*})}$ (or $|\textrm{supp}(p(z))|$), and hence a smaller $R$ and faster contraction. In contrast, a prior fully supported on $\{0,1\}^{p}$ results in an exponential dependence of $R$ on the dimensionality $p$, since $|\textrm{supp}\left(p(z)\right)|=2^{p}$. This analysis highlights the challenge of high-dimensional invariant prediction problems and the importance of an informative prior in that setting.
@@ 3.2 Finite-sample extensions and violations of the assumptions
First, consider finite samples. Suppose we observe a fixed set of environments and per-environment sample size $n\to\infty$. The consistency results and rates from Section 3.1 continue to hold, provided that $z^{*}$ remains identifiable within the observed environments. This condition requires sufficient heterogeneity among environments.
@@ 4 Synthetic study
We evaluate the BIP framework with simulated data under known distributions. With these simulations, we confirm our theory, explore the properties of uncertainty quantification, and empirically compare our approach to existing methods for invariant estimation. In what follows, BIP as a standalone method refers to the version with exact inference, and VI-BIP denotes its variational approximation. All experimental details are provided in Supplementary Section C. We summarize our findings as follows:
  • In § 4.1 we empirically verify the theory with $p=3$ features where exact inference for BIP is tractable. Our results confirm posterior consistency and contraction rates as the number of environments $E$ and per-environment sample size $n$ increase.
@@
  • In § 4.3 we compare BIP and VI-BIP to existing methods for invariant estimation. We study both a low-dimensional setting ($p=10)$ and high-dimensional setting ($p=450$). We show that BIP and VI-BIP recover the invariant feature selector in low dimensions; and for high-dimensions, VI-BIP is more scalable and accurate than existing methods, while BIP with exact inference becomes computationally infeasible.
@@ 4.3 Comparison to other methods
    Oracle: We fit a pooled linear regression on the true invariant features $x^{z^{*}}$ and return the ones significant at the $\alpha=0.05$ level. If $p+1>nE$, we use Lasso with 10-fold cross-validation and return the top 10 features by absolute coefficient values.
@@
    Regression: Similar to the oracle, but linear regression is performed on all features rather than restricting to the true invariant features.
@@
    EILLS (Fan et al., 2023): A linear regression model that introduces a regularization term to encourage invariant predictions among environments. Following the default choice in their codebase, the regularization penalty is set to 36.
@@
    Variational inference BIP (VI-BIP) [this paper]: Our method with variational inference (Algorithm 2). Again, we use the posterior mode as the VI-BIP estimate of the invariant selector. § B.2 provides more details about the optimization.
@@
We emphasize that VI-BIP does not require enumeration of candidate sets, and so more easily applies to high-dimensional settings. However, we found that the uniform prior over the full space of $\{0,1\}^{p}$ leads to unstable optimization when $p$ is large. In the high dimensional case ($p=450$), we set the prior to be uniform over $\{z\in\{0,1\}^{p}:\|z\|_{0}\leq 20\}$, i.e., a uniform prior over sets of 20 features. Note we still optimize over all sets of 20 features. This informative prior is significantly different from a screening step.
@@
Figure 3 (B) displays the posterior value at the true $z^{*}$ given by BIP. The posterior value converges to 1 with more environments. Additionally, the convergence is faster with higher intervention strength. This result is consistent with the theoretical contraction rates. We include the variational posterior value at $z^{*}$ from VI-BIP in Figure S3 in the supplementary material, where we observe a similar trend as in the exact inference, but with a slower convergence rate compared to the exact posterior.
@@
High-dimensional results.  Figure 4 gives the results for $p=450$. Panel A (Left) shows the exact discovery rate of the oracle and VI-BIP converges to 1 as $E$ increases. In contrast, the remaining methods exhibit a near-zero exact discovery rate across $E$, primarily due to failures in the pre-screening step.
Panel (A) also displays the coverage results. As expected, the gold-standard oracle maintains near-perfect coverage. The coverage of VI-BIP approaches 1 as more environments are observed. Notably, EILLS achieves high coverage at large $E$ by predicting an empty invariant feature set in these scenarios. The remaining methods have low coverage.
@@ 5 Gene perturbation study
Following Peters et al. (2016), this problem can be approached by invariant prediction. We treat the expression level of the target gene as outcome $y$ and the expression levels of the remaining $p=6,169$ genes as features $x$. The goal becomes to find a subset of feature genes $x^{z^{*}}\subset x$ that invariantly predict $y$ across environments. If an invariant feature $x^{\prime}\in x^{z^{*}}$ is perturbed, then the outcome $y$ is expected to change in a new environment created by this perturbation, which allows us to predict the effect of an gene prediction. For non-invariant feature $x^{\prime}\not\in x^{z^{*}}$, we do not make such prediction. Peters et al. (2016) also provides a justification through a causal perspective, where the invariant feature genes are viewed as the direct causes of the target gene.
@@ Appendix A Additional Theory and Proofs
We write $\forall_{p}a$ to mean “for $p(a)$-almost every realization of $a$” for a random variable $(a)$ a probability measure $p(a)$.
@@ A.5 Theory generalization under relaxed assumptions
We provide a generalization of Theorem 1 under more general conditions. This result supports the discussion in § 3.2 where some of the original assumptions are violated.
@@ Appendix C Synthetic data study
We first describe the generic data generative process, and then we provide details for each experiment in § C.1, C.2 and C.3.
@@
    Each conditional density is a linear Gaussian family, that is, $p(t^{(i)}\mid t^{(1:i-1)})=\mathcal{N}(t^{(i)}\mid\beta t^{(1:i-1)}+\beta_{0},\sigma^{2})$ for some $\beta,\beta_{0},\sigma^{2}$.
@@
    Observational environment $e=1$. For $i=1,\cdots,p+1$:
    • sample the intercept parameter from $\mathcal{N}(0,1)$,
    • sample the variance parameter from $\textrm{Uniform}\left([\sigma^{2}_{\min},\sigma^{2}_{\max}]\right)$, where $\sigma^{2}_{\min}$ and $\sigma^{2}_{\max}$ are lower and upper bounds on the variance parameter,
    • sample each coefficient value independently:
@@
      • if $\pi(i)\neq p+1$ i.e. $t^{(i)}=x^{(\pi(i))}$, (i) with probability $actprob\in[0,1]$ first sample its absolute value from $\textrm{Uniform}\left([lb,ub]\right)$ and then assign a random sign, (ii) otherwise set to 0.
@@
    Interventional Environments ($e=2,\cdots,E$) Based on the observational environment $e=1$, we first randomly draw a fraction $intfrac_{e}$ of features $x^{\prime}_{e}\subset x$ to be intervened on. Then, for $i=1,\cdots,p+1$:
    • If $t^{(i)}\not\in x^{\prime}_{e}$, set $p_{e}(t^{(i)}\mid t^{(1:i-1)})=p_{1}(t^{(i)}\mid t^{(1:i-1)})$
    • Otherwise, draw new parameters for $p_{e}(t^{(i)}\mid t^{(1:i-1)})$
      • sample intercept parameter by first sampling its absolute value from $\mathcal{N}(m_{e},1)$ and assigning to it a random sign, for some $m_{e}\in\mathbb{R}$ specific to the environment $e$.
      • sample variance parameter from $\textrm{Uniform}\left([\sigma^{2}_{e,\min},\sigma^{2}_{e,\max}]\right)$, where $\sigma^{2}_{e,\min},\sigma^{2}_{e,\max}$ are the lower and upper bounds of variance parameters specific to the environment $e$.
      • sample each coefficient value independently:
          with probability $changeprob_{e}$, copy the corresponding coefficient from $e=1$ case
          with probability $1-changeprob_{e}$, (i) with probability $actprob$ sample its absolute value from $\textrm{Uniform}\left([lb_{e},ub_{e}]\right)$, assign a random sign and (ii) otherwise set to 0.
@@ C.1 Empirical verification of theory
We include simulation details for the experiments in § 4.1. We set $p=3$, the maximum number of the true invariant features $p^{*}_{\max}=3$, and the minimum number $p^{*}_{\min}=1$. For $e=1$, the lower bound of the absolute coefficient value $lb=0.5$, upper bound $ub=2$; $actprob=1.0$; noise level lower bound $\sigma_{\min}=0.1$ and upper bound $\sigma_{\max}=0.2$. For $e=2,\cdots,E$,

### v2

@@ 1 Introduction
Originally, Peters et al. (2016) motivated these assumptions through a causal problem. Each environment arises by an unknown intervention on the features $x$; the outcome $y$ is directly caused by a subset of those features; and the goal is to analyze the multi-environment data to determine the direct causes of $y$. Here we solve the same problem but state the invariance assumption directly without appealing to causality. While these assumptions often arise from a causal problem, invariance has wider applications including prediction robustness (Bühlmann, 2020), domain adaptation (Rojas-Carulla et al., 2018), and out-of-distribution generalization (Arjovsky et al., 2019).
@@
The paper is organized as follows. In § 2 we state the invariance assumption and introduce BIP, including the model structure, the exact inference procedure, and a scalable variational approximation. In § 3, we present theoretical results on posterior consistency and contraction rates. In § 4, we validate BIP on simulations. In § 5, we use BIP to study real data. In § 6, we summarize BIP and discuss limitations and future work.
@@ 1.1 Related works
Since Peters et al. (2016), invariant prediction has become an active area of research. Heinze-Deml et al. (2018) extends this framework to non-linear settings. Rothenhäusler et al. (2019) proposes an invariance notion based on feature-residual inner products in linear models, linking it to causal inference under additive interventions. Fan et al. (2023) incorporates residual-based invariance as a regularization term in linear regression. Wang et al. (2024) establishes a theoretical connection between invariant prediction models and causal outcome models, leading to a computationally efficient algorithm for causal discovery. Gu et al. (2025) shows that solving for exact invariance is NP-hard and introduces an objective that interpolates between exact invariance and predictive performance. In concurrent research, Madaleno et al. (2025) leverages a Bayesian hierarchical model to explicitly test the invariance of causal mechanisms.
This paper offers a probabilistic perspective on invariant prediction, with goals closely aligned to Peters et al. (2016) and Fan et al. (2023). As we will see below, this perspective opens the door to both new algorithms and new theoretical understanding.
@@ 2.1 Multi-environment data and the invariance assumption
We call $z^{*}$ the (true) invariant feature selector and $x^{z^{*}}$ the (true) invariant features. By Assumption 1 and the chain rule, each environment’s joint distribution factorizes,
@@ 1 Initialize variational parameter $\phi_{0}$
In detail, we first estimate $\hat{g}(y\mid x^{z})$ and $\hat{p}_{e}(y\mid x^{z})$ from data, and then optimize Eq. 7 with stochastic optimization. Specifically, we use Monte Carlo approximations of gradients of the ELBO, a form of black box variational inference (Kingma and Welling, 2014; Ranganath et al., 2014). Algorithm 2 summarizes this procedure. The complexity of VI-BIP is $O(T\cdot M\cdot c(\mathcal{D},p))$ for $T$ iterations and $M$ gradient samples. When the cost of exact BIP inference is prohibitive, VI-BIP provides a good alternative.
A technical challenge in VI-BIP is computing gradients of the ELBO, which involve expectations with respect to the discrete variational distribution $q_{\phi}(z)$. Standard reparameterization gradients do not directly apply, as they require continuous latent variables. Here we use the unbiased U2G gradient estimator (Yin et al., 2020); see Algorithm 3 in the supplementary materials. For a comprehensive discussion of gradient estimation methods for discrete latent variables, see Mohamed et al. (2020). Additional implementation and optimization details for VI-BIP are provided in § B.2.
@@ 3 Theory
We use $\stackrel{{\scriptstyle P}}{{\to}}$ to denote convergence in probability, $\textrm{supp}\left(p(z)\right)$ for the support of $p(z)$, and $\forall_{p}z$ to mean for every $z$ with positive prior value. Additional results and proofs are included in Supplementary Section A.
@@ 3.1 Main results
$R$ $:=\left(\max_{z\neq z^{*}}\frac{p(z)}{p(z^{*})}\right)\cdot|\textrm{supp}\left(p(z)\right)|,$ (8)
@@
The first is the prior, which influences the term $R$ in Eq. 8. An informative prior that assigns lower probability to non-invariant $z$s will lead to smaller value of $\max_{z\neq z^{*}}\frac{p(z)}{p(z^{*})}$ (or $|\textrm{supp}\left(p(z)\right)|$), and hence a smaller $R$ and faster contraction. In contrast, a prior fully supported on $\{0,1\}^{p}$ results in an exponential dependence of $R$ on the dimensionality $p$, since $|\textrm{supp}\left(p(z)\right)|=2^{p}$. This analysis highlights the challenge of high-dimensional invariant prediction problems and the importance of an informative prior in that setting.
@@ 3.2 Finite-sample extensions and violations of the assumptions
First, consider finite samples. Suppose we observe a fixed set of environments and let per-environment sample size $n\to\infty$. The consistency results and rates from Section 3.1 continue to hold, provided that $z^{*}$ remains identifiable within the observed environments. This condition requires sufficient heterogeneity among the environments.
@@ 4 Synthetic study
We evaluate the BIP framework with simulated data under known distributions. With these simulations, we confirm our theory, explore the properties of uncertainty quantification, and empirically compare our approach to existing methods for invariant prediction. In what follows, BIP as a standalone method refers to the version with exact inference, and VI-BIP denotes its variational approximation. All experimental details are provided in Supplementary Section C. We summarize our findings as follows:
  • In § 4.1 we empirically verify the theory in § 3.1 with $p=3$ features where exact inference for BIP is tractable. Our results confirm posterior consistency and contraction rates as the number of environments $E$ and per-environment sample size $n$ increase.
@@
  • In § 4.3 we compare BIP and VI-BIP to existing methods for invariant prediction. We study both a low-dimensional setting ($p=10)$ and a high-dimensional setting ($p=450$). We show that BIP and VI-BIP recover the invariant feature selector in low dimensions; and for high-dimensions, VI-BIP is more scalable and accurate than existing methods, while BIP with exact inference becomes computationally infeasible.
@@ 4.3 Comparison to other methods
    Oracle: We fit a pooled linear regression on the true invariant features $x^{z^{*}}$ and return the ones significant at the $\alpha=0.05$ level.
@@
    Regression: Similar to Oracle, but linear regression is performed on all features rather than restricting to the true invariant features. If $p+1>nE$, we use Lasso (Tibshirani, 1996) with 10-fold cross-validation.
@@
    EILLS (Fan et al., 2023): A linear regression model that introduces a regularization term to encourage invariant predictions among environments. Following the default choice in their codebase, the invariance regularization strength $\gamma$ is set to 36.
@@
    Variational inference BIP (VI-BIP) [this paper]: Our method with variational inference (Algorithm 2). Again, we use the posterior mode as the VI-BIP estimate of the invariant selector.
@@
We emphasize that VI-BIP does not require enumeration of candidate sets, and so more easily applies to high-dimensional settings. However, we found that the uniform prior over the full space of $\{0,1\}^{p}$ leads to unstable optimization when $p$ is large. In the high dimensional case ($p=450$), we set the prior to be uniform over $\{z\in\{0,1\}^{p}:\|z\|_{0}\leq 20\}$, i.e., a uniform prior over sets of up to 20 features. Note we still optimize over all sets of 20 features. This informative prior is significantly different from a screening step.
@@
Figure 3 (B) displays the posterior value at the true $z^{*}$ given by BIP. The posterior value converges to 1 as the number of environments increases. Additionally, the convergence is faster with higher intervention strength. This result is consistent with the theoretical contraction rates. We include the variational posterior value at $z^{*}$ from VI-BIP in Figure S3 in the supplementary material, where we observe a similar trend as in the exact inference, but with a slower convergence rate compared to the exact posterior.
@@
High-dimensional results.  Figure 4 gives the results for $p=450$. Panel A (Left) shows that the exact discovery rate of Oracle and VI-BIP converges to 1 as the number of environments $E$ increases. In contrast, the remaining methods exhibit a near-zero exact discovery rate across $E$, primarily due to failures in the pre-screening step.
In Figure 4, Panel A (Right) displays the coverage results. As expected, Oracle maintains near-perfect coverage. The coverage of VI-BIP approaches 1 as more environments are observed. Notably, EILLS-s achieves high coverage at large $E$ by predicting an empty invariant feature set in these scenarios. The remaining methods have low coverage.
@@ 5 Gene perturbation study
Following Peters et al. (2016), this problem can be approached by invariant prediction. We treat the expression level of the target gene as outcome $y$ and the expression levels of the remaining $p=6,169$ genes as features $x$. The goal becomes to find a subset of feature genes $x^{z^{*}}\subset x$ that invariantly predict $y$ across environments. If an invariant feature $x^{\prime}\in x^{z^{*}}$ is deleted, then the outcome $y$ is expected to change in the resulting new environment, which allows us to predict the gene deletion effect. For non-invariant feature $x^{\prime}\not\in x^{z^{*}}$, we do not make such prediction. Peters et al. (2016) also provides a justification through a causal perspective, where the invariant feature genes are viewed as the direct causes of the target gene.
@@ Appendix A Additional Theory and Proofs
We write $\forall_{p}a$ to mean “for $p(a)$-almost every realization of $a$” where $a$ is a random variable $a$ and $p(a)$ is its probability measure.
@@ A.5 Theory generalization under relaxed assumptions
We provide a generalization of Theorem 1 under more general conditions. This result supports the second part of § 3.2 where some of the original assumptions are violated.
@@ Appendix C Synthetic data study
We first describe the general data generative process for the experiments in § 4.1 and 4.3, and then we provide details for each experiment in § C.1, C.2 and C.3.
@@
    Each conditional density is a linear Gaussian family, that is,
    for some parameters $\alpha_{ei}\in\mathbb{R}^{i-1},\beta_{ei},\sigma_{ei}\in\mathbb{R}$. For $i=1$ we set $p_{e}(t^{(1)}\mid t^{(1:0)})=\mathcal{N}(t^{(1)}\mid\beta_{e1},\sigma_{e1}^{2})$ with parameters $\beta_{e1},\sigma_{e1}\in\mathbb{R}$. (For convenience, $t^{(1:0)}$ denotes an empty vector.)
@@
    Observational environment ($e=1$). For $i=1,\cdots,p+1$:
    • sample the intercept parameter $\beta_{1i}$ from $\mathcal{N}(0,1)$,
    • sample the variance parameter $\sigma_{ei}^{2}$ from $\textrm{Uniform}\left([\sigma^{2}_{\min},\sigma^{2}_{\max}]\right)$, where $\sigma^{2}_{\min}$ and $\sigma^{2}_{\max}$ are lower and upper bounds on the variance parameter,
    • sample each dimension of the linear coefficient parameter $\alpha_{ei}$ independently:
@@
      • if $\pi(i)\neq p+1$ i.e. $t^{(i)}=x^{(\pi(i))}$, (i) with probability $p_{\text{act}}\in[0,1]$ first sample its absolute value from $\textrm{Uniform}\left([lb,ub]\right)$ and then assign a random sign, (ii) otherwise set to 0.
@@
    Interventional Environments ($e=2,\cdots,E$) For each interventional environment $e>1$, we first randomly draw a fraction $\rho^{\text{int}}_{e}$ of features $x^{\prime}_{e}\subset x$ to be intervened on – referred to as the intervention strength in the main text. Then, for $i=1,\cdots,p+1$, if $t^{(i)}\not\in x^{\prime}_{e}$, we set $p_{e}(t^{(i)}\mid t^{(1:i-1)})=p_{1}(t^{(i)}\mid t^{(1:i-1)})$; otherwise, we draw new parameters for $p_{e}(t^{(i)}\mid t^{(1:i-1)})$ as follows
    • sample the intercept parameter $\beta_{ei}$ by first sampling its absolute value from $\mathcal{N}(m_{e},1)$ and assigning to it a random sign, for some $m_{e}\in\mathbb{R}$ specific to the environment $e$.
    • sample the variance parameter $\sigma_{ei}^{2}$ by first sampling a multiplicative factor $\lambda\sim\textrm{Uniform}([\lambda_{e,\min},\lambda_{e,\min}+\lambda_{e,\textrm{diff}}])$, and then setting $\sigma_{ei}^{2}=\lambda^{2}\sigma_{1i}^{2}$, where $\lambda_{e,\min}$, $\lambda_{e,\textrm{diff}}$ are parameters specific to the environment $e$.
    • sample each dimension of the linear coefficient parameter $\alpha_{ei}$ independently:
      • with probability $p_{\text{change}}$, copy the corresponding coefficient from $e=1$ case
      • with probability $1-p_{\text{change}}$, (i) with probability $p_{\text{act}}$ sample its absolute value from $\textrm{Uniform}\left([lb_{e},ub_{e}]\right)$, assign a random sign and (ii) otherwise set to 0.
@@ C.1 Empirical verification of theory
We include simulation details for the experiments in § 4.1. We follow the general data generative process described in the preamble of Supplementary Section C with the following configuration:
We set $p=3$, $p^{*}_{\max}=3$, $p^{*}_{\min}=1$, and $p_{\text{act}}=1$.
For $e=1$, we set $lb=0.5$, $ub=2$; $p_{\text{act}}=1.0$, $\sigma_{\min}=0.1$ and $\sigma_{\max}=0.2$.
For $e=2,\cdots,E$, we sample $m_{e}\sim\textrm{Uniform}([0,1])$, $\lambda_{e,\min}\sim\textrm{Uniform}([0.1,0.2])$, and $\lambda_{e,\textrm{diff}}\sim\textrm{Uniform}([0.1,0.5])$, and we set $lb_{e}=(m_{e}+0.01)^{2}$, $ub_{e}=(m_{e}+0.5)^{2}$, and $p_{\text{change}}=1$.